OpenStudy (anonymous):

Functions f(x) and g(x) are shown below. f(x) = x^2 g(x) = x^2 − 8x + 16 In which direction and by how many units should f(x) be shifted to obtain g(x)? A. Left by 4 units B. Right by 4 units C. Left by 8 units D. Right by 8 units

OpenStudy (anonymous):

Ummm our fist step would be to convert g(x) into vertex form \( g(x)=x^2-8x+16\) \(g(x)=(x^2-8x+16)-16+16\) \(g(x)=(x-4)^2\)

OpenStudy (anonymous):

you know what might work google search it will help with many math problems http://www.wolframalpha.com/

OpenStudy (anonymous):

Now that our equation is in the form \(y=a(x-s)^2+t\) We know that our s represents how many units the graph has been shifted horizontally So if s is negative it has been moved to the right s units If s is positive it has been moved to the left s units Soo How many units has this graph been shifted?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

umm right or left?

OpenStudy (anonymous):

Left?

OpenStudy (anonymous):

Well is s positive or negative?

OpenStudy (anonymous):

Negative

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

correct :) Did you understand how I converted the equation to vertex form?